When asked, "Do kids regularly carry weapons in your school?" 14,250 students said yes, and 115,339 students said no. With a 95% level of confidence, find the margin of error (MOE) associated with the sample proportions. (Give the answer to one decimal place, if needed.)
Answer #1

Given : n=115339 , X=14250

The sample proportion is , p=X/n=14250/115339=0.1235

q=1-p=1-0.1235=0.8765

Now , ; From standard normal distribution table

Therefore , the margin of error (MOE) associated with the sample proportions ia ,